Marc Horowitz 21.512

Nerd Tropes

In the field of computer science, there is a rich vocabulary used to describe concepts and objects which are common in the hacker’s world. Examples of all kinds of tropes, including metaphors, metanyms, and synecdoche, can be found.

The most common kind of trope in computer science is the metaphor. There are many things which are unique to this field, but are easily compared to more common objects. For instance, a secure computer system may have a design which allows a user (sometimes legitimate, sometimes not) to bypass the security systems. The part of the design which makes this possible is called a “back door”. There are a few examples of body-related metaphors. In a large computer network, the main circuit or set of circuits over which signals travel is referred to as the “backbone” of the network, in the same way that a person’s backbone is the main pathway for signals traveling between the brain and the body. A system which is so poorly designed as to be nearly unusable will be described as “brain-damaged”. A technique in programming which is so obscure or complicated that few, if any, people understand it is usually called “magic”. Many of these tropes, while unusual-sounding at first, are common in hacker vocabulary. They provide a vehicle for communicating information which English does not have any other way of expressing.

Metanyms are most commonly found when referring to certain common books in the hacker literature. The technical name for a certain part of a common operating system is called a daemon (pronounced like “demon”). The standard reference book on this operating system has a picture of a “daemon” on the front. This book is often referred to as “the Devil Book”. A classic book on compiler design has a picture of a dragon, hence it is called “the Dragon Book”. There is a common graphics system which has a three-volume reference. Since red, green, and blue are the fundamental colors in graphics, the publisher cleverly gave the books red, green, and blue colors. Rather than referring to these books by their titles, hackers refer to them as “the Red Book”, “the Green Book”, and “the Blue Book”. These names are often used by “wizards” (someone who practices “magic”) since the titles are usually long and cumbersome.

There are also examples of synecdoche in the hacker jargon. One very common example is “suit”. Hackers have considerable disdain for managers, salesdroids, and other industry members who aren’t technically trained. Since you can identify them by their clothing (as compared to hackers, who will avoid suits at all costs), they are referred to as “suits”, since there is nothing else worth mentioning about them.

The various forms of tropes one can find in everyday speech among hackers demonstrates many of the values of their culture. Many words have positive or negative connotations or nuances of meaning which are made explicit by the trope used. But to an outsider (or a suit), these meanings may be difficult or impossible to distinguish. Or, they may be very clear.